OpenStudy (anonymous):

Which point is on the circle described by (x - 2)2 + (y + 3)2 = 4? answer gets a medal

OpenStudy (anonymous):

Which point are you referring to?

OpenStudy (anonymous):

like a point on a graph

OpenStudy (anonymous):

Use the center of the circle, given by your equation.

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

where the center of the circle is (h,k)

OpenStudy (shamil98):

-h, -k = - -2 , - + 3 = (2,-3)

OpenStudy (anonymous):

that's not even a answer choice